# MemLLM: Finetuning LLMs to Use Explicit Read-Write Memory

### v2

@@ Abstract
While current large language models (LLMs) perform well on many knowledge-related tasks, they are limited by relying on their parameters as an implicit storage mechanism. As a result, they struggle with memorizing rare events and with updating their memory as facts change over time. In addition, the uninterpretable nature of parametric memory makes it challenging to prevent hallucination. Model editing and augmenting LLMs with parameters specialized for memory are only partial solutions. In this paper, we introduce MemLLM, a novel method of enhancing LLMs by integrating a structured and explicit read-and-write memory module. MemLLM tackles the aforementioned challenges by enabling dynamic interaction with the memory and improving the LLM’s capabilities in using stored knowledge. Our experiments indicate that MemLLM enhances the LLM’s performance and interpretability, in language modeling in general and knowledge-intensive tasks in particular. We see MemLLM as an important step towards making LLMs more grounded and factual through memory augmentation.
@@ 1 Introduction
In this paper, we introduce MemLLM, an LLM endowed with an explicit memory component. It has the general advantages of some of the memory-focused work we discussed: we can keep information accessible indefinitely, beyond the context window, including infrequent information that standard LLMs struggle with.
The LLM has both read and write access to the memory component, i.e., it can store information in the memory as it processes text (or interacts with a user) and retrieve it when it needs it.
We specify an API for read and write access. The LLM issues read commands to retrieve from the memory and write commands to write to the memory. We adopt finetuning to train the model to access the memory. Based on the API specification, we create a dataset with training examples of API read and write commands and finetune the LLM on it. Our published training dataset can be used to finetune any language model, endowing it with explicit memory without requiring architectural changes.
Our memory has an explicit structured schema, similar to a database schema. It is interpretable and inspectable for humans. It is editable by humans. It is scalable since databases have excellent scalability properties. It is interoperable since the contents of the memory can be exported (e.g., to a different LLM supporting explicit memory) and contents can be imported from data resources (e.g., from Wikidata).
@@ 2 Related work
Our framework aligns with external memory methodologies but stands out with its structured format for storing information. This explicit memory facilitates large-scale knowledge editing and makes the model’s output generation process more interpretable. While similar structured storage approaches exist, they are often task-specific, such as data record management (Hu et al., 2023). Unlike these, our method is designed for generic language modeling, making it adaptable to a variety of tasks without extensive prompt engineering. Our published training dataset can be used to endow any trainable language model with explicit memory without requiring architectural changes.
@@ 3.1 Memory structure
where is the memory. That is, we retrieve all triples with entities/relations from the candidate sets such that their average similarity to query subject and relation is above the threshold. Subject queries are handled analogously.111We discuss how we set the thresholds and other hyperparameters in Appendix C.
@@ 3.3 Finetuning the LLM
The loss is applied to the API call (ii) – this teaches the model to generate the correct API call. The loss is also applied to the posttext – this teaches the model (a) to make good use of the information provided in the query result for predicting entities and (b) to predict the next memory read (as indicated by “({”). (iii) is not subject to the loss since the query results are generated by the memory, not by the LLM. For the training example that contains the very first “({MEM_READ(” in the document (and only for this ), the loss is also applied to the pretext – because the LLM needs to learn where to generate this first “({MEM_READ(”.
@@ 4.1 Setup
We finetune two Mistral-7B (Jiang et al., 2023) models using LoRA (Hu et al., 2022), a memory-write model and a memory-read model. See Appendix C for details on finetuning and hyperparameters. Our baselines are the original Mistral-7B and the memory-read model with its memory capabilities disabled. The latter baseline lets us ascertain to what extent improvements are due to in-domain finetuning (as opposed to the memory).
@@ 4.1.1 Filtering Distant Supervision Relations
We present the results in Table 1. These results suggest that the reasoning approach outperforms the other two approaches by a large margin. Also, it suggests that the filtering would lead to higher quality based on the high recall score, 0.84. We demonstrate the best-performing prompt in Appendix D.
@@ 4.2 Perplexity evaluation
Memory redundancy reduction benefits. Instead of storing facts in a structured triple format, each fact could be stored as its own proposition (Chen et al., 2024). For instance, one can store the triple as the proposition: “Washington D.C. is the capital of the United States.” Based on the memory populated with full Wikipedia (111M triples), an equivalent RAG system would need to store and index 111M proposition sentences. In a vector-based (dense) retrieval setting, this means storing and indexing 111M vectors alongside their text counterparts. However, in MemLLM, we store these facts in structured triples with identifiers (cf. Section 3.1), and the only vector-based indices are entity and relation embeddings. After storing the 111M triples, the entity/relation tables contain roughly 21M unique entity/relation records. Thus, the total number of vectors in MemLLM is less than one-fifth of what would be needed to index proposition sentences. Moreover, having only entities and relations encoded as vectors reduces ambiguity and improves recall by introducing less noise compared to encoding entire sentences.
@@ 4.3 Knowledge Editing Evaluation
Table 3 compares knowledge editing results for MemLLM with three baselines.444 Like other baselines, we do not compare RAGs in the knowledge editing task due to the challenge of modifying facts in their unstructured format. Editing would require altering every related instance in the KB, which is impractical. MemLLM’s memory is not pre-populated as a KB; it stores edited facts through a memory-write process, similar to non- and semi-parametric editing methods. MemLLM outperforms the baselines (AVG of .84). High reliability (.78) and generalization (.76) scores suggest that MemLLM (i) manages to extract and store the relation triple based on the edit and (ii) utilizes the edit in the memory-read process to answer the original and the generalization test questions correctly. Moreover, since MemLLM uses an explicit memory the applied edits only mildly affect the answers to unrelated questions: MemLLM has a score of .97 on locality. This indicates that there is little cumulative deterioration of the explicit memory.
@@ 5 Conclusion
While the structured relation-based memory improves factuality and interpretability, it has its own limitations. The current version of MemLLM handles only 96 relation types commonly used in Wikidata. However, to handle all types of knowledge extraction and storage, the model should be capable of extracting other types of relations. Composite relations that could be inferred from multiple already extracted relations are not detected or utilized in the current version of MemLLM. For instance, if we extract (California, country, United States) and (Apple Inc., located in, California), we expect the relation (Apple Inc., located in (or Country), United States) to be inferred. MemLLM is not a memory-aware solution. This means if a fact is not stored in the memory, but the decoding process generates a partial prompt that requires that fact, the model would either continue generation based on its parametric knowledge or hallucinate. We refer all these limitations to future work, as in this paper we have laid the initial groundwork for building a more complex and comprehensive method.
@@ Appendix B Filtering Ambiguous Queries
Appendix C Hyperparameters Details
@@
Appendix D Filtering Prompt

### v3

@@ Abstract
While current large language models (LLMs) perform well on many knowledge-related tasks, they are limited by relying on their parameters as an implicit storage mechanism. As a result, they struggle with memorizing rare events and with updating their memory as facts change over time. In addition, the uninterpretable nature of parametric memory makes it challenging to prevent hallucination. Model editing and augmenting LLMs with parameters specialized for memory are only partial solutions. In this paper, we introduce MemLLM, a novel method of enhancing LLMs by integrating a structured and explicit read-and-write memory module. MemLLM tackles the aforementioned challenges by enabling dynamic interaction with the memory and improving the LLM’s capabilities in using stored knowledge. Our experiments indicate that MemLLM enhances the LLM’s performance and interpretability, in language modeling in general and knowledge-intensive tasks in particular. We see MemLLM as an important step towards making LLMs more grounded and factual through memory augmentation. The project repository is publicly available at: https://github.com/amodaresi/MemLLM
@@ 1 Introduction
In this paper, we introduce MemLLM, an LLM endowed with an explicit memory component. This architecture has the following key characteristics.
This explicit memory component has the general advantages of some of the memory-focused work we discussed: we can keep information accessible indefinitely, beyond the context window, including infrequent information that standard LLMs struggle with.
The LLM has both read and write access to the explicit memory component, i.e., it can store information in the memory as it processes text (or interacts with a user) and retrieve it when it needs it.
We adopt finetuning to train the model to access the explicit memory component through read and write commands. To this end, we specify an API for read and write access. Based on the API specification, we create a dataset with training examples of API read and write commands and finetune the LLM on it. Our published training dataset can be used to finetune any language model, endowing it with an explicit memory component without requiring architectural changes.
The memory component has an explicit structured schema, similar to a database schema. Therefore, it is interpretable and inspectable for humans; it is editable by humans; it is scalable since databases have excellent scalability properties; and it is interoperable since the contents of the memory can be exported (e.g., to a different LLM supporting explicit memory) and contents can be imported from data resources (e.g., from Wikidata).
@@ 2 Related work
In general, our framework aligns with external memory methodologies but stands out with its structured format for storing information. This explicit memory facilitates large-scale knowledge editing and makes the model’s output generation process more interpretable. While similar structured storage approaches exist, they are often task-specific. For example, Hu et al. (2023) introduce ChatDB, which takes a database structure as input for a data record management task. This requires prior knowledge of the task-specific database schema (e.g., a sales table), which must be defined and provided as a prompt to the model. In contrast, our method is designed for generic language modeling, making it adaptable to a variety of tasks without extensive prompt engineering. Knowledge graphs are another structured format that can be interpreted as a memory of triples, similar to our memory. However, while there is work on extracting triples from text into knowledge graphs (Zhang & Soh, 2024) using LLMs and on LLMs “interactively” querying knowledge graphs to answer questions (Baek et al., 2023), our approach is distinguished by teaching the LLM memory-write and memory-read functionality through finetuning. This makes memory an integral component of the model’s text processing and at the same time enables the model to flexibly handle new and updated knowledge. Similar integrated systems have also been considered in the knowledge graph literature (e.g., AutoKG (Zhu et al., 2024)), but to the best of our knowledge, our system is the first that goes beyond a conceptual proposal and demonstrates empirical success in both language modeling and knowledge editing. Additionally, our published training dataset can be used to endow any trainable language model with explicit memory without requiring architectural changes.
@@ 3.1 Memory structure
where is the memory. That is, we retrieve all triples with entities/relations from the candidate sets such that their average similarity to query subject and relation is above the threshold. Subject queries are handled analogously.111We discuss how we set the thresholds and other hyperparameters in Appendix D.
@@ 3.3 Finetuning the LLM
The loss is applied to the API call (ii) – this teaches the model to generate the correct API call. The loss is also applied to the posttext – this teaches the model (a) to make good use of the information provided in the query result for predicting entities and (b) to predict the next memory read (as indicated by “({”). (iii) is not subject to the loss since the query results are generated by the memory, not by the LLM. For the training example that contains the very first “({MEM_READ(” in the document (and only for this ), the loss is also applied to the pretext – because the LLM needs to learn where to generate this first “({MEM_READ(”. For a more comprehensive overview of the memory-read data generation process, refer to Appendix C, which outlines the detailed algorithm.
@@ 4.1 Setup
We finetune two Mistral-7B (Jiang et al., 2023) models using LoRA (Hu et al., 2022), a memory-write model and a memory-read model. See Appendix D for details on finetuning and hyperparameters. Our baselines are the original Mistral-7B and the memory-read model with its memory capabilities disabled. The latter baseline lets us ascertain to what extent improvements are due to in-domain finetuning (as opposed to the memory).
@@ 4.1.1 Filtering Distant Supervision Relations
We present the results in Table 1. These results suggest that the reasoning approach outperforms the other two approaches by a large margin. Also, it suggests that the filtering would lead to higher quality based on the high recall score, 0.84. We demonstrate the best-performing prompt in Appendix E.
@@ 4.2 Perplexity evaluation
Memory redundancy reduction benefits. Instead of storing facts in a structured triple format, each fact could be stored as its own proposition (Chen et al., 2024). For instance, one can store the triple as the proposition: “Washington D.C. is the capital of the United States.” Based on the memory populated with full Wikipedia (111M triples), an equivalent RAG system would need to store and index 111M proposition sentences. In a vector-based (dense) retrieval setting, this means storing and indexing 111M vectors alongside their text counterparts. However, in MemLLM, we store these facts in structured triples with identifiers (cf. Section 3.1), and the only vector-based indices are entity and relation embeddings. After storing the 111M triples, the entity/relation tables contain roughly 21M unique entity/relation records. Thus, as the number of extracted facts increases, our structured triple approach results in significantly lower memory usage and, at large scales, it requires less than 20% of the memory (21M) needed for direct proposition storage (111M), demonstrating its efficiency in reducing redundancy (Figure 4). Moreover, having only entities and relations encoded as vectors reduces ambiguity and improves recall by introducing less noise compared to encoding entire sentences.
@@ 4.3 Knowledge Editing Evaluation
Table 3 compares knowledge editing results for MemLLM with three baselines. MemLLM outperforms the baselines (AVG of .84). High reliability (.78) and generalization (.76) scores suggest that MemLLM (i) manages to extract and store the relation triple based on the edit and (ii) utilizes the edit in the memory-read process to answer the original and the generalization test questions correctly. Moreover, since MemLLM uses an explicit memory the applied edits only mildly affect the answers to unrelated questions: MemLLM has a score of .97 on locality. This indicates that there is little cumulative deterioration of the explicit memory.
@@ 5 Conclusion
While the structured relation-based memory improves factuality and interpretability, it has its own limitations. The current version of MemLLM handles only 96 relation types commonly used in Wikidata. However, to handle all types of knowledge extraction and storage, the model should be capable of extracting other types of relations. Composite relations that could be inferred from multiple already extracted relations are not detected or utilized in the current version of MemLLM. For instance, if we extract (California, country, United States) and (Apple Inc., located in, California), we expect the relation (Apple Inc., located in (or Country), United States) to be inferred. MemLLM is not a memory-aware solution. This means if a fact is not stored in the memory, but the decoding process generates a partial prompt that requires that fact, the model would either continue generation based on its parametric knowledge or hallucinate.
We acknowledge that Retrieval-Augmented Generation (RAG) methods are widely used for a more grounded text generation. However, like other baselines in knowledge editing, we do not include RAG methods in our comparisons due to the difficulty of modifying facts in their unstructured format. Updating a single fact would require locating and altering every related text snippet and embedding within the RAG’s knowledge base, which is highly impractical. MemLLM’s memory is not pre-populated as a KB; it stores edited facts through a memory-write process, similar to non- and semi-parametric editing methods. While methods like ChatDB (Hu et al., 2023) also employ non-parametric structured memory, they rely on database tables that are more suitable for analytical tasks. This setup requires the user to define a task-specific database schema in advance and explicitly prompt the model with it. In contrast, MemLLM uses a much more generalized memory structure that supports both language modeling and QA tasks (as discussed in the knowledge editing section) without requiring any changes to its approach or memory format. Furthermore, the nature of prompt-based approaches—compared to fine-tuning—makes them inherently less faithful and more dependent on the specific content stored in memory.
We refer all these limitations to future work, as in this paper we have laid the initial groundwork for building a more complex and comprehensive method.
@@ Appendix B Filtering Ambiguous Queries
Appendix C Memory-read Data Generation
Algorithm 1 presents the pseudocode for the process of generating MemLLM ’s memory-read training data. See Section 3.3 for a detailed description of the same process.
Appendix D Hyperparameters Details
@@
Appendix E Filtering Prompt